### v1

# BlendScape: Enabling Unified and Personalized Video-Conferencing Environments through Generative AI

Abstract.
Today’s video-conferencing tools support a rich range of professional and social activities, but their generic, grid-based environments cannot be easily adapted to meet the varying needs of distributed collaborators. To enable end-user customization, we developed BlendScape, a system for meeting participants to compose video-conferencing environments tailored to their collaboration context by leveraging AI image generation techniques. BlendScape supports flexible representations of task spaces by blending users’ physical or virtual backgrounds into unified environments and implements multimodal interaction techniques to steer the generation. Through an evaluation with 15 end-users, we investigated their customization preferences for work and social scenarios. Participants could rapidly express their design intentions with BlendScape and envisioned using the system to structure collaboration in future meetings, but experienced challenges with preventing distracting elements. We implement scenarios to demonstrate BlendScape ’s expressiveness in supporting distributed collaboration techniques from prior work and propose composition techniques to improve the quality of environments.
Overview figure of the key components, findings, and visuals results of BlendScape with the following components: (1) A high-level view of the BlendScape user interface, showcasing a user entering a prompt, a generated environment with composited user video feeds, and a control row for environment adjustments; (2) A summary of key system features: (a) Different composition modes, (b) automatic layout techniques based on object detection, and (c) techniques for adding custom objects and artifacts; (3) A summary of the findings from our evaluation (a) preserving backgrounds enables personal and natural environments, (b) spatial richness vs. cohesiveness of composition, (c) varied preferences for strong vs. subtle theming, (d) context-informed layouts for co-presence, (e) Full generation for social context, subtle composition for professional settings, (f) Potential distractions and required time to generate; (4) A selection of example environments generated by BlendScape.
1. Introduction
Distributed collaboration tools are becoming essential to the way we conduct daily activities. Since the pandemic, there has been a rise in popularity of using video-conferencing tools for a rich set of collaboration scenarios, ranging from professional tasks (e.g., remote work and distance learning (Benabdallah et al., 2021; Lee et al., 2022)) to social gatherings (e.g., connecting with family members (Fuchsberger et al., 2021), practicing music (Cai et al., 2021), playing tabletop games (Yuan et al., 2021)). However, today’s video-conferencing tools do not reflect this wide variety of activities due to how these tools compose meeting environments (i.e., the “stage” or background rendered around users’ videos). Users are typically placed in separate regions of a video grid within generic rooms that lack a connection to the meeting context, which can lead to meeting fatigue (Fauville et al., 2021), reduce user engagement (Choi and Diehm, 2021), and disrupt interpersonal cues necessary for mediating conversations (Tang et al., 2023; Hyrkas et al., 2023; Hu et al., 2023a).
In recent video-mediated communication and HCI systems research, we observe three approaches to reimagining the composition of video-conferencing environments to better support distributed collaboration: (1) Establishing a shared context, e.g., through rendering users’ videos in a unified environment (O’Hara et al., 2011; Morikawa and Maesako, 1998) and setting the mood or theme of a meeting (Hunter et al., 2014; Follmer et al., 2010); (2) Enabling spatial metaphors for communication (e.g., by incorporating proxemic interactions (Hu et al., 2023a) and spatial sound (Hyrkas et al., 2023; Tang et al., 2023)) to help users mediate turn-taking); (3) Using the environment to record a meeting history, to provide a basis for later collaboration (e.g., through immersive authoring capabilities (Xia et al., 2018)). However, it is still a challenge for end-users to adapt meeting environments to meet their distributed collaboration needs during live meetings (Grønbæk et al., 2023a). Customization tools in commercial systems (e.g., Microsoft Teams Together Mode111Microsoft Teams TogetherMode: https://www.microsoft.com/en-us/microsoft-teams/teams-together-mode, Ohyay222Ohyay: https://ohyay.co/) require significant time and effort, which may lead users to default to generic, pre-defined options.
In this paper, we present BlendScape, a rendering and composition system that enables video-conference participants to easily create environments tailored to the meeting context, with two key innovations: (1) We ground the generation of meeting environments in real spaces that are meaningful to users by blending their physical or virtual backgrounds into a unified environment. This can serve as a mechanism for personalization (Venolia et al., 2018) or to incorporate physical objects to collaborate around (Hu et al., 2023b; Junuzovic et al., 2012). (2) Capitalizing on recent advances in generative AI, we leverage image generation models to enable expressive and rapid techniques for composing environment designs. While such techniques are the subject of ongoing research, several dominant modes of use have emerged:
text-to-image: An image is generated from a given text prompt.
image-to-image: An image is generated from a text prompt and an input image, retaining features of the input image while introducing new elements or styles consistent with the prompt.
inpainting: Similar to image-to-image, but where an image mask is used to determine which parts of the input image should be unchanged. The rest of the image is generated in a way that it is consistent with the fixed parts of the input image (i.e., blended).
BlendScape uses (1) an inpainting technique that creates a new environment by merging users’ webcam backgrounds into a unified environment; (2) an image-to-image technique that transforms an existing image of an environment (i.e., an image prior) to reflect the meeting purpose and composites users’ videos within it. We enable meeting participants to steer the generation of environments through multimodal interaction techniques, including text-based prompts, direct manipulation, and choice of image priors. To lower the barrier for end-users to generate good quality meeting spaces, we developed composition techniques in BlendScape to guide the AI models to produce contextually relevant environments and integrate users’ videos within the scene in an immersive manner.
We evaluated BlendScape in two steps. First, to demonstrate the expressiveness of BlendScape’s composition techniques (Ledo et al., 2018), we implemented and analyzed three scenarios using BlendScape. These scenarios, guided by a review of prior work, incorporate a range of seven techniques for supporting distributed collaboration (e.g., conveying physical and digital task spaces, enabling ad-hoc layouts of users that reflect the meeting context, or recording a history of collaboration within a meeting space). Our analysis focused on the extent to which BlendScape can be used to implement these techniques in familiar or new ways.
Second, we conducted a user study with 15 end-users to investigate their preferences for customizing environments for various meeting scenarios and to what extent BlendScape enables them to achieve their design intentions. Study participants compared and critiqued pre-generated environments for two social gatherings, then used BlendScape to design a series of meeting spaces for a remote work scenario. The participants found value in BlendScape’s composition techniques for enabling them to rapidly prototype environment designs for a wide range of meeting activities. However, they experienced challenges with mitigating distracting and unrealistic elements to achieve the ideal design. We propose extensions and improvements to BlendScape’s implementation to address these limitations in future work.
Our work contributes: (1) the BlendScape system, which enables end-user customization of video-conferencing environments through generative AI-driven composition techniques; (2) three scenarios that demonstrate how BlendScape can be used to prototype a range of seven techniques for supporting distributed collaboration from prior work; (3) an evaluation with 15 frequent users of video-conferencing tools, which surfaced benefits and limitations of BlendScape for creating customized meeting spaces.
2. Related Work
Our work builds on prior research on (1) composing video-conferencing environments to support distributed collaboration; (2) generative AI techniques for constructing 2D and 3D environments.
2.1. Video-Conferencing Environments
To situate our work in the video-mediated communication literature, we first classified techniques for composing meeting environments, based on a review of commercial video-conferencing tools and HCI systems research. Here, we define the environment as the “stage” or background rendered around users’ videos, excluding other functional elements of video-conferencing tools such as the chat or host controls. As shown in Fig. 2, we further categorized these composition techniques according to three overarching benefits they can provide for distributed collaboration, as demonstrated through empirical studies in prior work: (1) establishing a shared context, (2) enabling spatial metaphors for communication, (3) serving as a record or artifact of collaboration.
Classification of Environment Composition Techniques. This figure shows a table with three columns. The first column has header ”Establishing a Shared Context” and children ”Rendering users’ videos within a unified environment”, ”Conveying the physical or digital task space”, ”Setting the theme or mood of a meeting. The second column as header ”Enabling Spatial Metaphors for Communication” and children ”Enabling proxemic metaphors”, ”Enabling scaling metaphors”, ”Rendering spatialized audio”, ”Establishing consistent user layouts”, ”Rendering unique viewing perspectives.” The third column has header ”Serving as a Record or Artifact of Collaboration” and children ”Recording a history of collaborative activities”, ”Serving as an artifact produced through collaboration.”
We focus our review on screen-based meeting systems with 2D or 2.5D designs; composition techniques for other modalities (e.g., tabletop (Zagermann et al., 2016; Tuddenham and Robinson, 2009) or mixed reality interfaces (Pejsa et al., 2016; Herskovitz et al., 2022; Grønbæk et al., 2023b)) would likely introduce new dimensions to our categorization.
Establishing a Shared Context.
A primary role of virtual meeting environments is creating a shared frame of reference for distributed collaborators. One approach popularized by commercial tools (e.g., Teams Together Mode, Ohyay, Zoom Immersive View333Zoom Immersive View: https://blog.zoom.us/introducing-zoom-immersive-view/) as well as research systems (e.g., Waazam (Hunter et al., 2014), HyperMirror (Morikawa and Maesako, 1998), BISi (O’Hara et al., 2011)) is rendering users’ videos within a single, unified environment to simulate a co-located meeting. Prior studies found this approach promising for enhancing user engagement and increasing co-presence, as compared to traditional grid-based video layouts (Tang et al., 2023; Hunter et al., 2014). Teams and Zoom provide lightweight tools for uploading custom environment images and manually placing participants, whereas Ohyay offers an extensive customization suite for meeting organizers to design multi-room environments and script interactive behaviors. Yet, the process of tailoring environments to align with the meeting context can be highly manual and time-consuming with these tools, which may cause users to default to pre-defined templates.
Prior meeting tools have explored a variety of environment design techniques for conveying the task space, i.e., the physical or digital space where work takes place and artifacts are collaboratively produced (Buxton, 2009). Commercial tools like Teams and Zoom enable screen-sharing capabilities to display digital task spaces (e.g., slideshows, shared documents), which are rendered alongside users’ videos in a grid-based layout. MirrorBlender (Grønbæk et al., 2021) enables blending the person and task space through “mirrors” (i.e., translucent representations of users’ video feeds and shared screens) that can be layered on top of each other to enable more natural interaction techniques for hybrid collaboration (e.g., using hand gestures to interact with shared content). Conveying physical task spaces (e.g., for learning hands-on skills (Kumaravel et al., 2019) or practicing music (Cai et al., 2021)) is an ongoing challenge with commercial video-conferencing tools, often requiring users to implement custom hardware setups to capture their physical surroundings (Junuzovic et al., 2012; Labrie et al., 2022). To mitigate this, ThingShare (Hu et al., 2023b) and VideoPlay (Follmer et al., 2010) enable users to scan physical objects and manipulate them as digital representations during video calls.
Enabling Spatial Metaphors for Communication.
To enable more natural and seamless communication, recent tools’ meeting environments are designed to leverage spatial affordances that mimic social cues and collaboration strategies from face-to-face interactions. Our review surfaced five spatial composition techniques (Fig. 2): enabling interpersonal interactions through manipulating proximity and scale of users’ videos, rendering spatial audio, structuring users’ layouts to support turn-taking,
A common approach is using proxemic metaphors to facilitate conversations: in Gather444Gather: https://www.gather.town/, users can initiate one-on-one video calls by “walking up” to each other in a 2D map; OpenMic (Hu et al., 2023a) enables users to approach a “Virtual Floor” in the center of the environment to express their desire to speak. Scaling metaphors are commonly used to highlight specific users: Microsoft Teams and Zoom offer Speaker views that render active speakers at a larger size than other meeting participants; OpenMic (Hu et al., 2023a) allows users on the Virtual Floor to negotiate conversational transitions by increasing the size of their videos. To help users follow the flow of conversations, Hyrkas et al. explored requirements for rendering spatialized audio for co-located and remote users in hybrid meetings (Hyrkas et al., 2023); MirrorVerse (Grønbæk et al., 2023a) implements a “doorway” metaphor that enable users to preview conversations in breakout rooms through spatial audio. Systems like Teams TogetherMode and Ohyay place users in unified 2D environments that establish consistent user layouts (e.g., by seating users around a table), which can help establish turn-taking patterns. Perspectives (Tang et al., 2023) further supports turn-taking by rendering unique viewpoints for distributed users, which can enable more natural behaviors from co-located meetings, e.g., turning your head and making eye contact with active speakers.
Serving as a Record or Artifact of Collaboration.
We observed a few examples where meeting environments can document collaborative activities, providing a basis for later iterations on ideas.
First, meeting environments can record a history of user layouts and changes to the task space, which can be replayed at a later time to understand collaboration patterns. For example, MirrorVerse (Grønbæk et al., 2023a) offers record-and-replay tools to capture changes in the entire workspace, including how users reconfigure video layouts and progressively adapt meeting controls to their needs. Finally, the meeting environment could be an artifact produced through collaborative activities, e.g., for interior design or world-building usage scenarios. In our review of prior work, we did not find examples of video-conferencing tools that explicitly claim this functionality; however, this use case is common in VR immersive authoring tools (Xia et al., 2018; Zhang et al., 2023a) which allow users to collaboratively create 3D environments while being immersed in the environment itself.
Our review of prior work surfaces a wide range of environment composition techniques that can enhance users’ awareness, communication, and sense of co-presence in distributed meetings. However, it is still a challenge for meeting participants to tailor environments to enable these affordances as needed, due to the time and manual effort required to compose environments with existing customization suites (Grønbæk et al., 2023a). With BlendScape, we contribute a flexible set of generative AI-enabled techniques that can be used to create and adapt meeting environments to serve ad-hoc collaboration activities. In Sec. 5, we return to this classification of prior video-conferencing systems to demonstrate the range of distributed collaboration techniques that BlendScape can enable.
2.2. AI-Assisted Environment Generation
Generative AI techniques have been a focal point of prior HCI research (Agarwala et al., 2004; El-Nouby et al., 2019; Dang et al., 2022; Liu et al., 2022; Dang et al., 2023), which explored the interplay between generative models and user-driven design interactions. Based on user input, such as visuals or textual prompts, these models can generate a wide variety of content including images (Avidan and Shamir, 2007; Mirza and Osindero, 2014; Zhang et al., 2017; Yu et al., 2019; Rombach et al., 2022), text (Radford et al., 2018; Brown et al., 2020; OpenAI, 2023), and even 3D objects (Coyne and Sproat, 2001; Gkioxari et al., 2020; Koh et al., 2022; Karnewar et al., 2023; Jun and Nichol, 2023).
Recent advances in image generation techniques allow the generation to be conditioned in a variety of useful ways, in addition to guidance from text-based prompts. For example, ControlNet uses an auxiliary model that incorporates information conveyed through additional input modalities, such as depth, semantic, and human pose representations (Zhang et al., 2023b; Li et al., 2023). Prior work also explored intuitive techniques for end-users to control generative AI models, e.g., through sketches (Chung et al., 2022; Zhang et al., 2023b), voice (Numan et al., 2023), painting (Chung and Adar, 2023b), continuous and discrete sliders (Chung and Adar, 2023a) and iterative design mechanisms (Sharma et al., 2018; El-Nouby et al., 2019; Dang et al., 2023).
There is increased interest in harnessing image generation models for the dynamic generation of virtual environments (Coyne and Sproat, 2001; Dang et al., 2023). Before the rise of AI-driven approaches, environment generation relied on predefined components and procedural logic. For instance, WordsEye converts text descriptions into corresponding 3D scenes using a fixed database of 3D models and predefined rules (Coyne and Sproat, 2001). Oasis transforms 3D-scanned physical spaces into procedurally generated virtual environments, mapping walkable areas and recognizing real-world objects to enhance immersion (Sra et al., 2018). DreamWalker used procedural techniques to substitute real-world elements with virtual ones, enabling users to stay immersed in VR while safely navigating the real world (Yang et al., 2019).
These systems paved the way for the utilization of generative techniques that can be observed in recent AI-assisted creativity support tools. A recent example in this domain is WorldSmith, a system that investigates how multimodal image generation systems can be harnessed to aid users in visualizing and iteratively refining elements of fictional worlds (Dang et al., 2023). Opal (Liu et al., 2022) guides users through a structured search for visual concepts to generate images for news illustrations, utilizing GPT-3 for prompt generation including keyword, tone, icon, and style suggestions based on an article’s content.
BlendScape builds on similar techniques observed in these systems, but with a distinct focus on supporting distributed collaboration. While WorldSmith employs inpainting methods for enhancing elements in fictional 2D worlds, BlendScape utilizes inpainting to blend users’ backgrounds into unified and contextually relevant meeting environments. Like Opal, BlendScape leverages a large language model (LLM) to dynamically generate prompts aligned with user-specified meeting themes and activities.
Overview of ”Product Design Ideation” scenario with the following components: (a) Blank canvas with original user video feeds. (b) Blended environment with user video feeds composited into the scene based on prompt ”brainstorming”, with region-based user-defined prompt specifying ”wall with post-it notes” for granular environment editing. (c) Blended environment with addition of wall with post-it notes according to user input. (d) Blended environment with user video feeds composited into the scene based on prompt ”design studio with the design of sketches of prototypes”.
3. Scenario Walkthrough
To illustrate how BlendScape can enable expressive meeting environments by blending physical and digital spaces, we present an initial scenario implemented with our system: a brainstorming session between two interaction designers who are collaborating to design a new mixed reality interface.
The designers join a video call and add their webcam feeds to BlendScape’s composition interface; one designer also adds a camera feed of their physical notebook in their design studio, so they can sketch ideas during the brainstorming session (Fig. 3a). The designers resize and reposition their videos through click-and-drag interactions, to render the whiteboard large and in the center of the environment. They type “brainstorming” in the prompt field to generate a creative environment for their planned design activity. A few seconds later, they see their blended environment: part of their physical surroundings are still visible, but they now extend into one unified design lab (Fig. 3b) where their own video feeds seamlessly blend in. This blended design allows them to gesture and refer to physical sketches in the notebook, simulating how they might interact in a co-located design studio. To preview mock-ups of their MR interface, the designers blend a digital sketch into their physical backgrounds (Fig. 3c).
Many other meeting experiences are possible (Appendix. A.2), from work scenarios to therapy sessions, birthday parties, vacation or wedding planning. As we describe in the next section, BlendScape provides flexible and modular techniques for creating such rich meeting spaces. Later, we discuss additional scenarios that demonstrate the expressiveness of BlendScape to enable distributed collaboration techniques from prior work (Sec. 5).
4. BlendScape System
Overview of BlendScape’s user interface: (a) Two composition modes – inpainting using a webcam feed and image-to-image techniques for refining canvas images, (b) Automatic layout techniques for compositing users behind objects and scaling videos, (c) Input fields for prompts specifying Meeting Activity and Meeting Theme, (d) Slider for users to specify the extent of video background preservation, (e) Strength of Stylization slider to control the priority of stylistic prompts or user backgrounds, (f) Transform tool for moving and scaling user video feeds, (g) Selection tool to mark area for region-based modification, and (h) Button to upload an image to the canvas.
We present BlendScape, a system for meeting participants to compose and customize video-conferencing environments. Key to our approach is blending elements of users’ physical or virtual backgrounds into unified environments, to incorporate visuals that are meaningful to the collaboration context or meeting participants. We first present three system requirements that guided the design of BlendScape. Then, we introduce a set of generative AI techniques for creating blended environments and composition techniques for enhancing the visual cohesion and naturally integrating users within the scene.
4.1. Requirements
We defined three requirements for generating blended environments with BlendScape, informed by our motivating scenario and our review of video-conferencing literature:
R1: Enabling embedding meeting context into the environment. Our Design Brainstorming scenario (Sec. 3) and review of environment composition techniques for distributed collaboration (Sec. 2.1) surfaced many examples where the environment is used to reflect the collaboration context, e.g., conveying a shared task space (Grønbæk et al., 2021) or establishing a meeting theme (Hunter et al., 2014). BlendScape enables this through an image inpainting technique, which generates new environments by blending users’ webcam backgrounds together, and an image-to-image technique, which transforms input images of overarching environments and composites users within it. We also implemented prompt enhancement techniques that utilize LLMs to encourage the generation of functional and stylistic elements to support specific meeting scenarios.
R2: Compositing users in a cohesive and immersive manner. To create the illusion of meeting in a shared space, we aimed to render users’ videos such that they appear naturally integrated with the environment. One key strategy is preserving a portion of users’ video backgrounds while generating blended spaces. This makes the region around the user match the physical conditions and noise present in their input video stream (i.e., maintaining realistic lighting, shadows, and color temperature). We also developed hidden surface removal techniques in BlendScape to place users behind foreground objects, obscuring the boundaries of their webcams.
R3: Enabling coarse- and fine-grained customization of environments. Prior video-conferencing tools demonstrated a need for meeting participants to update the entire scene (e.g., generating spatial layouts to support small vs. large group discussions (Grønbæk et al., 2023a)) as well as make granular adjustments (e.g., adding objects as visual aids (Hu et al., 2023b)). In addition to BlendScape’s coarse-grained inpainting and image-to-image techniques for generating new environments, we enable fine-grained adjustments by selecting portions of the scene to re-generate. We also maintain a history of generated environments to enable iteration on past results.
4.2. BlendScape Interface
Next, we provide an overview of the BlendScape system (Fig. 4). BlendScape’s user interface consists of a canvas in the center that displays blended environments and composites the users’ videos within them. Meeting participants can generate environments via two composition modes (Fig. 4a): leveraging the video feed from only their webcam (using inpainting), or refining the image that is already present in the canvas (using image-to-image techniques). This image on the canvas may be a result of a previous image generation step or an image uploaded by the user (Fig. 4f). Users can specify to what extent their video backgrounds should be preserved through a slider (Fig. 4i).
BlendScape allows users to enter two prompts for a Meeting Activity and Meeting Theme (Fig. 4c) to generate visuals relevant to the purpose of their meeting. To control whether the prompts or the users’ backgrounds should be prioritized in the environment generation, users can change the Prompt Strength (Fig. 4b). BlendScape offers direct manipulation techniques (via clicking, dragging, and pinching) as intuitive ways to position and scale users, both for steering the environment generation and placing users’ videos in the resulting scene.
To make fine-grained adjustments, users can add or remove objects by selecting a region of the environment (Fig. 4e) and specifying a prompt. BlendScape also offers automatic layout techniques (Fig. 4d) that composite users behind objects in the scene and automatically scale their videos to match the size of the objects. BlendScape saves a history of past environment generations and users’ positions within the scene, to enable iterating on past designs (Fig. 4g).
4.3. Generative AI Techniques for Blended Environments
BlendScape implements two classes of AI image generation techniques for composing blended video-conferencing environments: (1) an inpainting technique, which blends users’ physical or virtual backgrounds into a unified environment, and (2) an image-to-image technique, which incorporates users into an overarching environment that represents the meeting setting.
Generating blended environments with users’ video backgrounds (R1, R2)
With BlendScape’s inpainting techniques, meeting participants can blend their real-world surroundings or virtual backgrounds into the shared environment as a mechanism for personalization (Venolia et al., 2018) or capturing the task space where artifacts are collaboratively produced (Buxton, 2009). First, users can specify the proportion of their video backgrounds to retain; we mask these regions to preserve them in the blended environment (Fig. 6). Then, BlendScape performs inpainting to generate plausible visual details between the fixed regions, based on user-specified prompts of Meeting Activity and Meeting Theme, e.g., ‘‘brainstorming’’ in a ‘‘treehouse’’-themed environment (Fig. 5a). Preserving physical backgrounds, rather than generating fully artificial scenes, makes users appear more naturally integrated into the blended space, as their real-world conditions (e.g., lighting, shadows, color temperature, and webcam resolution) are reflected in the resulting scene.
To steer the generation of different styles of environments, users can reposition and scale their videos, e.g., positioning smaller images at the top of the screen can create a sense of depth, with a clear separation of foreground and background elements.
An overview of environment generation techniques in a flow diagram. BlendScape supports composing meeting spaces through (a) blending video feeds together via inpainting techniques and (b) transforming an input image (i.e., image prior) via image-to-image techniques. These composition approaches can be chained, e.g., to restyle a blended environment in the theme of a library (c).
Overview of techniques for masking video backgrounds. Via BlendScape’s video background controls (Fig. 4), users can adjust the proportion of their physical or virtual surroundings to mask or retain in the resulting blended environments. Four images are shown in a 2-by-2 grid. In the top-left image, the original video backgrounds are shown on a blank canvas. In the other three corners, different inpainted canvases are shown. These canvases are generated with different levels of background preservation (minimum, medium, maximum), where minimum results in the least amount of influence of the user’s background video on the generated canvas and maximum refers to the most influence on the generated canvas.
Driving environment generation through image priors (R1)
As a second composition approach, BlendScape uses image-to-image techniques to transform an image prior (i.e., input image). This involves restyling the environment to incorporate visuals related to the Meeting Activity and Meeting Theme prompts, while preserving key structural elements in the prior, e.g., walls or background objects that convey the environment’s geometry (Fig. 5b).
Meeting participants can upload image priors to set the theme of the meeting, (e.g., using an image of a library for a study group meeting) or establish a layout of users (e.g., seating in rows for a college lecture). Furthermore, the image-to-image technique could be a strategy for re-creating a real-world meeting space that is familiar to distributed collaborators.
Combining inpainting and image-to-image techniques for iterative environment design. (R1, R3)
To simplify novice users’ composition process into two clear pathways for creating environments, we separated the inpainting and image-to-image techniques into two modes within BlendScape: webcam-based and canvas-based generation. However, BlendScape supports flexibly combining these techniques to iterate on prior environments.
We found that performing an image-to-image transformation after an inpainting generation is an effective strategy for refining roughly-blended areas and subtly restyling an existing environment (Fig. 5c), through prototyping usage scenarios (Sec. 5) and observing participants’ design process in our User Study (Sec. 6). BlendScape also enables blending users’ video backgrounds into an image prior, to incorporate authentic elements of their backgrounds while adhering to an overarching environment structure.
4.4. Improving the Composition of Blended Environments
Next, we describe three composition techniques to improve the quality and make granular adjustments to the generated environments: (1) prompt-enhancement techniques to provide more detailed instructions to image generation models; (2) granular controls for scene-editing to add visuals or correct distortions; (3) hidden surface removal techniques to integrate users’ videos with foreground objects.
Generating context-informed and thematic environments through LLM-driven prompt padding (R2)
Crafting expressive prompts remains a challenge for novice users of image generation models, requiring significant trial and error (Brade et al., 2023; Chung and Adar, 2023b). To lower the barrier to entry for meeting participants to create high-quality and detailed environments, BlendScape enhances user-specified prompts with additional keywords suggested by large language models, a technique demonstrated in prior systems research (e.g., Opal (Liu et al., 2022)). We elicit text-based prompts from users in two steps to decouple requirements for the layout of the environment, which is driven by the Meeting Activity, and aesthetic elements that reflect the Meeting Theme. Then, query GPT-3.5 with the Meeting Activity and Meeting Theme prompts as input, instructing the model to return keywords for five relevant objects in the meeting space and five stylistic qualities that represent the meeting atmosphere. The advantage of this approach is dynamically eliciting keywords that are tailored to the meeting context while minimizing the effort and expertise required from meeting participants. For example, for a Meeting Activity of “Brainstorming Session” and Meeting Theme of “Hologram,” GPT-3.5 suggested “Interactive Touchscreens, Holographic Whiteboards Floating Desks as objects and “Dynamic Lighting Effects, Seamless Integration of Virtual and Physical Elements, Collaborative and Hi-tech Atmosphere” as stylistic qualities. Inspired by prompt expansion tools for artists (e.g., PromptGen555PromptGen: https://promptgen.vercel.app/), BlendScape also adds a fixed set of terms to encourage high-quality results: “highly detailed, intricate, sharp focus, smooth.” Appendix A.2 includes full examples of GPT queries and outputs.
Compositing users in an immersive manner through hidden surface removal (R2)
A limitation of capturing meeting participants via webcams is that they appear as “floating heads” with a harsh cut-off at their shoulders, due to the limited field-of-view. BlendScape takes a similar approach to address this issue as prior tools (e.g., Ohyay, TogetherMode): placing users behind objects in the scene and enabling hidden surface removal, a rendering technique that removes portions of 3D objects that should not be visible from a particular camera perspective. We implemented an object segmentation pipeline in BlendScape to compute the salient objects in the generated environment (e.g., chairs, tables) and extract them to a foreground layer positioned in front of users’ videos (Fig. 9). This creates the illusion of users sitting behind objects, hiding the bottom edge of their videos.
Making granular edits to the environment (R3)
In addition to BlendScape’s coarse-grained techniques for updating the entire video-conferencing environment, we implemented a selection technique for making fine-grained revisions to the scene (Fig. 7). Using the selection tool, users can circle areas of the scene and specify a text-based prompt to remove content (e.g., to fix distorted areas) or add visuals relevant to the meeting scenario (e.g., additional chairs to accommodate new meeting participants). To regenerate the area, BlendScape uses the GLIGEN (Li et al., 2023) inpainting model, which is trained to consider the prompt, position, and scale of the specified region in relation to its surroundings, to generate a more cohesive and realistic result.
Region-based modification of blended backgrounds, with two examples. Firstly, a ”maker space” blended environment is modified by adding a whiteboard based on a user-defined prompt and region. Secondly, a ”library” blended environment is modified by adding a bookshelf based on a user-defined prompt and region.
BlendScape Components and Architecture. The BlendScape system consists of a Unity client that serves as the main user interface. Via Microsoft Teams NDI capabilities, the Unity client receives and processes users’ videos by performing person segmentation to separate the users from their video backgrounds and using inpainting techniques to fill the missing areas in the backgrounds. The Unity client connects to two servers that run (1) Stable Diffusion image generation processes to enable the environment generation techniques; (2) PixelLib for object segmentation to generate foreground objects, GPT-3.5 to enhance users’ prompts with contextually relevant keywords, and GLIGEN (grounded text-to-image generation model) for re-generating small portions of the environment.
4.5. Implementation
Figure 8 shows BlendScape’s key components and system architecture. We implemented BlendScape’s user interface as a Unity application integrated with Microsoft Teams. BlendScape receives video streams from individual meeting participants via Microsoft Teams’ NDI streaming capabilities.
Image generation models: To enable the environment generation techniques, we used open-source Stable Diffusion (Rombach et al., 2022) and ControlNet (Zhang et al., 2023b) models through the WebUI API 666Stable Diffusion WebUI: https://github.com/AUTOMATIC1111/stable-diffusion-webui, ControlNet Extension: https://github.com/Mikubill/sd-webui-controlnet, hosted on a PC with an AMD EPYC 7742 64-Core Processor and NVIDIA RTX A6000 GPU. For the inpainting technique, we used the Realistic Vision 2.0 777https://huggingface.co/SG161222/Realistic_Vision_V2.0 Stable Diffusion checkpoint which is finetuned for inpainting, further guided by a ControlNet inpainting model. For the image-to-image technique, we used the Realistic Vision’s base checkpoint, guided by ControlNet Depth and Canny models to preserve the spatial layout and salient features of the image priors.
Furthermore, we incorporated GLIGEN (Li et al., 2023) to allow users to regenerate specific areas of the scene. Similar to ControlNet (Zhang et al., 2023b) in its goal of guiding diffusion models, GLIGEN makes region-specific edits based on textual prompts and bounding box coordinates.
Segmentation of the generated scene: BlendScape uses PixelLib (Olafenwa, 2021) to perform semantic segmentation to partition the scene into meaningful segments across the depth and for hidden surface removal. To segment users from their video backgrounds, we used conventional computer vision techniques. While BlendScape displays live video feeds of the segmented users, we generate static environments using the first frame of users’ videos, due to limitations with computing Stable Diffusion outputs at video rates. This results in an empty space in the video backgrounds when we perform person segmentation, since the users’ webcams do not capture areas that the user is occluding. We used Stable Diffusion to inpaint this area to simulate a continuous background.
2.5D, layered scenes: The inpainting and image-to-image techniques require isolating different components of the scene (i.e., image-to-image transforms the entire existing environment while inpainting only takes the users’ webcam backgrounds as input). To assemble the correct input for the image generation models, we implemented BlendScape as a 2.5D scene in Unity with five 2D layers staggered at different depths: (1) foreground objects, (2) users’ videos (isolated from their video backgrounds), (3) users’ video backgrounds, (4) background masks (i.e., to preserve regions of the video backgrounds), and (5) the environment (as shown in Fig. 9). We instrument the Unity scene with multiple orthographic cameras that only render specific layers, enabling us to capture each layer separately. Since orthographic cameras do not perform perspective rendering, scene elements are rendered at a consistent scale even when placed at different depths.
Decomposition of the BlendScape scene to demonstrate the layered approach. Separate elements of the scene are grouped into layers and subsequently rendered with orthographic cameras, enabling the combination of possibly multiple environment generation passes. Layers include (from back to front): blended background image, person layer, foreground layer.
5. Demonstration of Additional Scenarios
In the first part of our summative evaluation, we use scenario demonstration (Ledo et al., 2018) as a method for illustrating the range of examples that can be achieved with our system. Our review of prior work surfaced ten environment composition techniques for supporting distributed collaboration (Sec. 2.1), e.g., enabling proxemic metaphors, incorporating representations of the task space, or recording a history of collaborative activities. In this section, we describe how we used BlendScape (Ledo et al., 2018) to implement three distributed collaboration scenarios (Design Brainstorming, Remote Education, Storytime with Family) that incorporate seven of these techniques. Comparing our generated meeting spaces to prior video-conferencing environments, we analyze how BlendScape can be used to implement these techniques similarly to prior work or introduce new composition strategies.
5.1. Design Brainstorming
Recall our initial scenario from Sec. 3, where two interaction designers are collaborating remotely (Fig. 3). They use BlendScape to render their videos in a unified workspace that blends together their physical and digital task spaces, allowing them to refer to hand-drawn and digital sketches.
Implementation with BlendScape: We prototyped this scenario using three live video feeds: two videos of the designers captured through laptop webcams, and one external camera feed capturing a physical desk with notebooks and a digital sketch (Fig. 3b).
Comparison to Prior Work: Similar to MirrorBlender (Grønbæk et al., 2021), BlendScape environments afford layering users’ videos on top of shared content to enable more natural interactions with the task space (e.g., using deictic expressions, pointing to content via hand gestures). However, BlendScape approach of blending video feeds together into a seamless depiction of the physical and digital task spaces moves beyond “windowing” environments in Teams, Zoom, or video conferencing research systems (e.g.,(Grønbæk et al., 2021)), where elements of the task space are displayed in separate blocks. While further research is required to understand the impact of these blended spaces on collaboration workflows, prior empirical studies of video-conferencing (Tang et al., 2023) and mixed reality collaboration systems (Pejsa et al., 2016; Grønbæk et al., 2023b) suggest that depicting a spatial arrangement of users within a unified space can increase co-presence and improve conversation flows.
5.2. Remote Education
Next, to demonstrate how BlendScape environments could support proxemic and spatial metaphors to facilitate conversations (Hu et al., 2023a), we prototyped a scenario involving a remote university course.
Scenario Description: A professor uses BlendScape to facilitate activities in their literature seminar course. To establish the room layout for small group discussions, they upload an image prior of a room with several tables and restyle it as a “library” (Fig. 10a, c). Students split into groups by organizing themselves around the tables, leveraging proxemic metaphors (Fig. 10c). To structure a larger group discussion afterwards, the professor restyles an image of a conference room (Fig. 10b, d). To spotlight students who are speaking, they use scaling metaphors by rendering the students’ videos larger and at the head of the table.
Overview of ”Remote Education” scenario with the following components and flow. (a, b) Image priors of a wedding venue and conference room. (b) Restyled environment with the wedding venue prior, using a prompt of ”library.” Chairs and tables are highlighted in blue to indicate objects that users can place themselves behind. (c) Restyled environment with the conference room prior, using a prompt of ”library”.
Implementation with BlendScape: We used BlendScape’s image-to-image mode with a Meeting Theme prompt of “library” to transform the image priors. To prototype the proxemic and scaling metaphors, we used BlendScape’s direct manipulation techniques.
Comparison to Prior Work: In contrast with OpenMic (Hu et al., 2023a), which leverages abstract spatial metaphors to enable proxemic interactions (i.e., a Virtual Floor boundary), BlendScape’s environments introduce literal and concrete spatial landmarks for users to organize around (e.g., conference room table with chairs). This is similar to Gather’s approach, which uses walls to indicate meeting space boundaries and podiums to designate areas to broadcast audio. However, a key difference from Gather is that BlendScape’s environments are not pre-canned; unexpected spatial properties could encourage new collaborative behaviors (e.g., collaborators could be compelled to split into groups if the room geometry introduces barriers, as shown in Fig. 13). Future work is needed to study the impact of abstract vs. concrete and deterministic vs. stochastic spatial properties on distributed collaborators’ behavior.
5.3. Storytime with Family
Our final scenario is inspired by systems like Waazam (Hunter et al., 2014) and VideoPlay (Follmer et al., 2010), which mediate playful social interactions between children and family members.
Scenario Description: A grandmother uses BlendScape to craft a memorable storytelling experience with her granddaughter. First, with a prompt of “storytelling,” a castle appears in the blended space between their video backgrounds (Fig. 11b). As the grandmother reads a fairy tale, she restyles the scenes to set the theme of the meeting; the environment transforms into magic ballroom and then a mushroom forest (Fig. 11c, e). She and her granddaughter reposition their videos in the scene to enact the roles of different characters. A few weeks later, the granddaughter wants to write a sequel to the fairy tale. They use BlendScape’s saved environments and video positions to replay a history of their collaboration and iterate on their previous scenes to author a new story.
Overview of ”Storytelling” scenario with the following components and flow. (a) Original video feeds. (b) Blended environment with user video feeds composited into the scene based on prompt ”storytelling”. (c) Restyling of blended ”storytelling” environment based on prompt ”magic castle, ballroom”. (d) New blended environment with user video feeds composited into the scene based on prompt ”mushroom forest”. (e) Iterative regeneration of blended ”mushroom forest” environment.
Implementation with BlendScape: We used the inpainting mode to blend the users’ physical environments together. We used image-to-image techniques with prompts of “magic castle, ballroom” and “mushroom forest” to restyle the blended environment to reflect the progression of the story.
Comparison to Prior Work: Existing environment customization tools (e.g., offered by Waazam (Hunter et al., 2014), VideoPlay (Follmer et al., 2010), or Teams TogetherMode) allows users to set the theme by choosing from preset backgrounds or curating and uploading their own images. This results in a manual and time-consuming customization process, but gives end-users fine-grained control over the resulting environment. On the other hand, BlendScape’s image-to-image restyling techniques offer a more rapid and automated method for creating themed environments, which could be more feasible for changing environments to support different collaborative activities during live meetings. BlendScape also enables making granular edits to scenes using GLIGEN inpainting (Li et al., 2023), however, end-users have less control over the output due to the generative model’s stochasticity.
Our review of prior video-conferencing systems (Sec. 2.1) surfaced few examples of using the environment to record a meeting history (e.g., MirrorVerse’s workspace recording tools (Grønbæk et al., 2023a)). BlendScape’s environment history could enable new approaches for facilitating distributed meetings, e.g., capturing visuals that represent discussion topics on the walls of the meeting room and holding later meetings in the same environment to kickstart follow-up conversations.
6. User Study
In the second part of our summative evaluation, we investigated how end-users could envision using BlendScape to personalize video-conferencing environments for various meeting scenarios. We conducted a user study with 15 individuals who regularly use traditional video-conferencing tools (e.g., Microsoft Teams, Zoom), where they used BlendScape to generate video-conferencing environments for various meeting scenarios. Our goals were to (1) Surface end-users’ preferences for customizing meeting environments, including visual or layout characteristics they aimed to achieve; (2) Investigate the benefits and challenges of BlendScape for allowing end-users to express their design intentions.
The study was approved by the Institutional Review Board within our organization. We conducted 1-hour study sessions remotely via Microsoft Teams. Participants were compensated with a $50 gift card.
6.1. Participants & Recruitment
We recruited individuals from our organization who use video conferencing tools at least three times per week, advertising through internal mailing lists. 15 individuals participated in our study (7 women, 8 men, majority in an age range of 25-34 years) and held a variety of job roles: UX or Product Designer (5), UX Researcher (5), Product Manager (3), Software Engineer (1), and Human Factors Engineer (1). 14 of 15 participants were located in North America and one was located in Asia.
6.2. Method
The study consisted of three environment composition tasks using BlendScape and a post-task discussion. In the first two tasks, we gradually taught participants to use BlendScape’s composition techniques and asked them to assess the quality of example environments in a semi-structured interview portion. Tasks 1 and 2 served as training for Task 3, where the participants composed a series of video-conferencing environments for a multi-step scenario.
Set-up: We hosted BlendScape on a local machine and gave participants remote control via Microsoft Teams screen-sharing to complete the tasks. For all tasks, we used pre-recorded videos to represent different users (simulated as NDI streams via the OBS Studio NDI Integration Tools888OBS Studio NDI Integration Tools: https://obsproject.com/forum/resources/obs-ndi-newtek-ndi™-integration-into-obs-studio.528/), in order to achieve relatively consistent environment generations across participants. We used a combination of real physical environments and virtual backgrounds to represent a diverse range of environments. For Task 3, we integrated participants’ webcam feeds into BlendScape so they could experience being immersed in different environments.
Task 1: Walkthrough & Comparison of Inpainting Techniques (10 min). First, we explored blended environment designs for a Vacation Planning scenario, where two friends are planning a trip to Paris and generate meeting spaces to reflect landmarks they want to visit. We chose this scenario to teach participants how even basic input (i.e., providing prompts, changing the position and scale of the image priors) can be used to steer the environment generation via BlendScape’s inpainting techniques. After the facilitator demonstrated a few prompts, we asked participants to experiment with different prompts and layouts.
We then showed participants three environments that preserved the users’ webcam backgrounds to varying degrees (Fig. 15). We asked participants to comment on which examples, if any, they could envision using for the Vacation Planning scenario with minimal changes. For the environments they could not envision using, we asked them to explain 1-2 key issues.
Task 2: Walkthrough & Comparison of Image-to-Image Techniques (10 min). Next, we introduced BlendScape’s canvas-based generation techniques using a Game Stream scenario involving a Minecraft player who uses BlendScape to make their viewers feel more engaged and immersed in the gameplay. We demonstrated the image-to-image approach, starting from an image of an arcade and using prompts to restyle the image prior in a Minecraft theme. Then, we demonstrated a more complex approach where we blended the users’ webcam backgrounds into the arcade image. Similarly to the first task, we asked participants to compare four example environments (Appendix A.3), comment on which options they could envision using for the scenario, and explain key issues they observed. We intentionally included examples of roughly blended and cluttered environments to provoke discussions around the limitations of BlendScape’s composition techniques.
Task 3: Designing Environments for a Progressive Meeting Scenario (15 min). To explore how participants would adapt environment designs during a live meeting, we instructed them to compose a series of environments for a scenario involving writing a research paper. First, we introduced a student character (using a pre-recorded video) and brought the participants’ webcam feeds into the prototype, to play the role of another student.
We facilitated the task via the following prompts: (1) Design an initial environment for two students to discuss the introduction of the paper; (2) The professor joins the call to provide feedback on the students’ ideas. How would you redesign the environment to include them? (3) All three users are starting to feel stressed with the paper deadline approaching. How would you redesign the environment to support them?
Throughout the task, we asked the participants to think aloud to describe their design goals and assess the quality of the generated environments. We prompted them to use features that we had not yet explored (e.g., adding or removing content).
Discussion (15 min). We ended the study with a semi-structured interview around the benefits and limitations that blended representations of environments could provide to remote collaboration scenarios. We asked participants to comment on specific scenarios where they could or would not envision using a system like BlendScape to create personalized and unified environments. Participants also reflected on unexpected or surprising elements of the environments they generated in the previous tasks and how these elements could support or detract from collaborative processes.
6.3. Data Collection & Analysis
For all study sessions, we collected audio transcripts & recordings, screen recordings, and images of the environments generated with BlendScape for later analysis. To analyze the environment characteristics that participants’ aimed to achieve, we used a thematic analysis approach (Braun and Clarke, 2006). One author reviewed all transcripts to create an initial codebook, specifically paying attention to participants’ rationale for selecting an environment over another in the Task 1 & 2 comparison exercises and revision strategies in Task 3. Three authors then reviewed the codes in the context of specific examples provided by the participants (including both quotes and images of the environments) to extract higher-level themes. We used an affinity diagramming approach (Scupin, 1997) to analyze and aggregate themes from the post-task discussions around the benefits and limitations of BlendScape and future usage scenarios.
Overview of participant environment designs for the Research Paper Scenario. To support the students’ and professor’s writing process in our Task 3 scenario, participants adopted a variety of design strategies: blending their backgrounds into spaces that support creativity (e.g., P6’s coffee shop and P8’s office space), creating environments themed according to the research topic (e.g., P2’s chemistry lab), and removing content to create empty, distraction-free spaces (P13). The participants envisioned both playful and calming environments to “de-stress” the users as they approached their paper deadline, e.g., P4’s rainforest with hammocks, P12’s stuffed animal-themed room, and P2 & P15’s indoor spaces with plants and natural lighting.
7. Results
In this section, we first present four themes from our analysis of participants’ environment customization preferences (i.e., the visual and layout characteristics participants aimed to achieve in their designs). We then discuss two themes around the benefits of BlendScape’s expressive composition techniques and limitations with the time and effort required to achieve optimal designs. Figure 12 shows a sample of the environments generated by the 15 participants in our study.
7.1. Environment Customization Preferences
We discuss four themes around participants’ preferences for (1) authentic over artificial spaces; (2) strongly vs. subtly themed environments; (3) structuring collaboration through spatial layouts of users; (4) balancing the spatial-richness of environments.
Authentic environments were preferred over artificial, but came with higher expectations for realism.
A majority of participants expressed that blending users’ physical surroundings into a unified meeting space can promote co-presence while maintaining familiar aspects of their individual environments (P2, P5-6, P9, P11-14). P13 commented that “taking something personal to [them] and tweaking it” would help remote collaborators to “trick [their] minds to believe that [they’re] more in the same place.”
P6 cited another benefit of authentic backgrounds in creating a higher degree of realism, as users’ videos appeared to be naturally framed with appropriate furniture and lighting. In environments constructed from virtual backgrounds, users sometimes appeared to be “floating in space in front of an image… like a bad Photoshop job” (P6). However, we observed a drawback to this increased sense of realism: participants more easily identified and were more critical of flaws in how their physical surroundings were blended (P1, P3-5, P7-8, P10), e.g., warped areas or inconsistent room geometry (Fig. 13). With fully artificial environments, some participants argued that “there’s no pressure to do it well” (P8); “even if it was not as polished” they would have a “higher tolerance” for mistakes (P3).
Varied preferences for strong vs. subtle visual ties to the meeting context.
A few participants appreciated having stylistic elements that closely reflect the Meeting Activity and Theme prompts they provided, e.g., a chemistry lab for the Research Paper scenario in Task 3. Despite some unrealistic visuals, P15 appreciated the departure from traditional meeting spaces that are “not really imaginative,” “fixed and static”. P7 commented: “it’s nice that we don’t need the space to look like traditional meeting spaces.” However, most participants preferred simple and subtle theming to avoid distractions (P1, P7, P9, P13-15); P1 expressed environments should “add texture to the call without pulling away from it.”
Environments that establish spatial layouts of users can help structure collaboration, but require manual staging.
A popular design strategy was placing users around spatial landmarks in the scene to reflect the collaboration context (P1, P3-7, P9, P13-15). For example, in the Research Paper scenario, P14 positioned the professor in the lower left corner “to help structure the collaboration,” similar to a picture-in-picture style for online lectures. P9 placed users in chairs that were close together to “give [them] the most privacy” while having a one-on-one discussion.
However, some participants observed that BlendScape did not always consider the existing layout of users when generating meeting spaces. P1 and P6 had to manually reposition users to seat them in chairs, but they expected BlendScape to automatically generate chairs and tables to frame the users. This points to a limitation of how BlendScape blends environments after segmenting users from their video backgrounds, which we further discuss in Sec. 8.2.
Balancing the spatial richness of environments with 2D representations of users.
Spatial properties of BlendScape’s environments (e.g., furniture layouts, lighting, and shadows aligned with the geometry of the space) increased realism and lent themselves to structuring user layouts for collaboration (as discussed in the previous theme). However, many participants observed a juxtaposition between these spatially-rich environments and the strictly 2-dimensional representation of users (P5-6, P8-10, P11, P13-14). A few participants tried to tilt the users’ videos to “echo the [3D] perspective in the room” (P14) and make it look like users were facing each other (P9, P13). To make our hidden surface removal technique (i.e., hiding floating heads behind foreground objects) more convincing, P14 suggested to “accept the flatness” and use environments where all furniture and users’ videos are facing forwards, in the style of Teams TogetherMode.
Two visual examples of imperfections in blended environments. Firstly, an environment with warped room geometry due to the disparate video backgrounds of the users. Secondly, an environment with warped room geometry due to the different perspectives of the users’ video backgrounds.
7.2. Benefits and Limitations of Composition Techniques
We discuss two additional themes around participants’ perceptions of the benefits and challenges with using BlendScape to compose meeting environments.
BlendScape provides expressive controls for creating environments that reflect the meeting context.
The participants appreciated BlendScape’s expressive power and ability to rapidly generate and iterate on environment designs with the system, creating over 300 environments across all three study tasks (115 for Task 1, 65 for Task 2, 125 for Task 3). As P13 expressed, they exerted “such minimal effort” with the system, and “[BlendScape ] does so much” to translate the prompts into a rich environment.
We also observed participants using a variety of design strategies that utilized the full range of our multimodal techniques to steer image generation. To prototype environments for the Research Paper scenario, they crafted prompts to generate settings that spark creativity (e.g., libraries, parks, coffee shops) (P1, P5-9, P11-12, P15), added objects that represent the users’ research topic (P2-3), and removed content from the environment to promote distraction-free collaboration (P10, P13, P14). To help “de-stress” the students as their deadline approached, a popular strategy was crafting meeting spaces that transport the users to relaxing or playful locations, e.g., a beach, a rainforest, and a stuffed animal-themed space (Fig. 12). Participants also used the image-to-image mode to subtly restyle the existing scene (e.g., using prompts of “warm” and “relaxing”).
Potential challenges with distracting scene elements, time & effort required to achieve the ideal design.
To enable using a system like BlendScape for professional contexts, participants expressed a need to prevent unexpected elements “that could prove to be more distracting than helpful” (P6). These ranged from minor mistakes that “look good when you first glance at it” (P5) but become more apparent upon inspecting the scene (e.g., Two Eiffel Towers in the Task 1 example environments as shown in Fig. 15) to more significant cases where the generative AI models misunderstood participants’ intent (e.g., GLIGEN inserting a playground slide rather than presentation slides for P9).
While each environment generation takes minimal effort, participants anticipated these issues requiring several iterations to fix (P1, P3-5, P7, P10, P12, P15). As P12 stated, “If we get it right, it will be good. If we get it wrong, then we’ll spend time trying to get it right.” To prevent disrupting meetings, some participants would prefer to generate personalized environments before, rather than during, a meeting (P1-3, P5).
8. Discussion
In our Discussion, we first reflect on opportunities to support distributed collaboration with BlendScape. Then, we review two challenges with BlendScape’s composition techniques and describe how to improve upon our implementation in future work. We end by discussing the limitations of our research.
8.1. Opportunities for Supporting Distributed Collaboration with BlendScape
We developed BlendScape to enable tailoring meeting environments to a wide range of distributed collaboration scenarios. All participants expressed they could envision using a system like BlendScape to generate video-conferencing environments in the future. In addition to professional contexts, for which participants preferred subtler environments that minimize distractions, they saw potential in BlendScape to support social scenarios (e.g., team-building activities, calls between friends and family members, board game nights, or baby showers).
Another goal of BlendScape was to support prior systems’ techniques to increase co-presence and improve conversation flows. It was encouraging that study participants, when using BlendScape for the first time, recognized potential collaboration affordances that blended environments could offer and prototyped examples similar to prior work (e.g., adding visuals to represent a shared task space (Hunter et al., 2014; Follmer et al., 2010), organizing users around spatial landmarks to support turn-taking in conversations (Hu et al., 2023a; Tang et al., 2023)).
While our evaluations focused on BlendScape’s expressiveness and user experience considerations, further research is needed to evaluate the impact of blended spaces on collaboration behaviors during live meetings. We outline three directions for future work, aligned with our classification of prior systems (Sec. 2.1):
(1) Capture and composition techniques for establishing a shared physical context. While prototyping the Design Brainstorming scenario, we observed that current inpainting models produced more realistic results when the capture perspective and scale of the physical task space match those of typical BlendScape environments (e.g.,, a whiteboard captured from landscape perspective). The overhead view of the physical desk often resulted in the task space being merged with the floor. Future work could study environment composition approaches for more fine-grained physical tasks that need to be recorded from a close-up view or birds-eye perspective (e.g., soldering, knitting).
(2) Abstract vs. concrete spatial metaphors for communication. BlendScape differs from other tools leveraging proxemic interactions (e.g., MirrorBlender (Grønbæk et al., 2021), OpenMic (Hu et al., 2023a)) by offering literal representations of spatial landmarks for users to organize around (e.g., chairs for sitting, walls to designate boundaries). Future comparative studies should investigate how users utilize abstract vs. detailed spaces differently and the impact on conversation flow. In particular, we anticipate challenges when users’ real-world surroundings are blended, as some of our study participants expressed discomfort at placing their videos in other users’ “personal space”.
(3) System-driven techniques for updating environments and recording collaboration history. The current BlendScape implementation saves a history of manual customizations that users make to the meeting space. As LLMs and image generation techniques advance towards operating at video rates, there could be an opportunity for a more automated approach to adapting environments to the meeting context. For example, through processing transcripts of the meeting (Chen et al., 2023), the system could infer an ideal environment design for the current meeting activity and load prior environments for discussing recurring topics.
8.2. Improving Upon BlendScape’s Composition Techniques
We developed BlendScape to achieve a balance between usability requirements for video-conferencing and technical constraints of current image generation models. Since it is infeasible to run current models at video rates, we rely on static backgrounds. However, our customization techniques (based on text-based prompts and direct manipulation) still enable users to easily and rapidly update environments to reflect the meeting progression. Our strategy of blending video streams together still allows for rendering dynamic portions of the scene by incorporating external camera feeds (as demonstrated in our Design Brainstorming scenario for conveying a physical workspace).
While we expect model performance and output quality to improve in the future, we propose extensions to BlendScape to improve upon two key limitations that were surfaced in our study.
(1) BlendScape can produce environments that are agnostic to the number and existing placement of users in the scene. Our current BlendScape implementation segments users from their backgrounds before performing inpainting to create blended environments. This was an intentional choice to avoid generating inappropriate or distorted depictions of humans, which is a known issue with image generation models. However, this creates a tradeoff that environments may not reflect the existing spatial layout of users in the scene (e.g., containing an overabundance of furniture for the number of users).
To achieve more “people-informed” compositions, we envision using OpenPose models999ControlNet OpenPose: https://huggingface.co/lllyasviel/sd-controlnet-openpose, which perform human pose estimation, to generate furniture layouts around existing users in the scene (e.g., automatically framing them with chairs and tables). This approach would likely require multiple passes: first, an inpainting step to generate the environment, using an input image with users present to support the OpenPose model. A second inpainting pass could be performed to remove any distorted depictions of people.
(2) The environment generation techniques can be overwhelmed when users’ backgrounds and text-based prompts are in contrast. Our study participants noticed that when users’ “backgrounds are so disparate, [BlendScape ] really had trouble integrating them into a believable environment” (P13). We observed this in particular with the environments in Task 2 (Fig. 16), where the four video streams had semantically different backgrounds (e.g., medieval castle vs. office space), with little relation to the prompt of Game Streaming.
To achieve more cohesive blended environments, we envision “prioritizing” or preserving a larger proportion of video backgrounds that resemble the meeting prompts (e.g., by computing similarity scores between the meeting prompts and semantic descriptions of users’ video backgrounds).
8.3. Limitations
We end our Discussion by reflecting on the limitations of our work.
Study Sample: First, we evaluated BlendScape with individuals who primarily work in user experience design and product management. Our participants’ opinions may not generalize to all potential users of BlendScape, in particular end-users outside of a design background. However, our participants had a variety of prior experience with generative AI: 7 of 15 participants use image generation models once a week or more frequently, and the remaining 8 use them once a month or not at all. This suggests that BlendScape’s composition techniques could be accessible to novice or first-time end-users.
Lack of Baseline Comparison: Given that we designed BlendScape to support a variety of meeting scenarios and a combination of distributed collaboration approaches from prior work, no singular meeting system serves as an effective baseline to compare the full range of BlendScape’s composition techniques against. Instead, we adopted a more holistic evaluation approach to demonstrate what BlendScape can enable: (1) implementing and analyzing usage scenarios, which serves as an initial comparison of BlendScape to systems from prior work; (2) studying usability considerations and customization preferences specific to BlendScape.
Ecological Validity: We pre-recorded videos with a variety of static backgrounds to simulate video calls in our study. This was to keep the environment generations relatively consistent across participants, so we could align and compare their feedback in our analysis. However, using primarily physical and dynamic video backgrounds could have elicited different design strategies and challenges from our participants (e.g., preferences towards periodically updating meeting environments to reflect changes in users’ surroundings). In future work, we plan to conduct studies using BlendScape in live meetings to investigate its affordances for supporting distributed collaborative behaviors.
Another limitation stems from using primarily virtual backgrounds in our study. Most participants chose not to use their physical environments in Task 3 (for privacy reasons or due to their surroundings being uninteresting). Although most participants did not have a first-hand experience of blending their real-world surroundings into the environment, we find it encouraging that they still brainstormed a variety of scenarios where using authentic environments could provide value, e.g., to connect family members or encourage co-workers to “share their lives in different ways” (P9).
Novelty Effects: Considering the novelty of generative AI techniques, our studies around BlendScape are subject to participant response bias (Dell et al., 2012). We reduced potential bias by discussing both the benefits and limitations of BlendScape with participants and probing them with examples of poorly designed environments (e.g., some of the Game Stream environments shown in Fig. 16) to elicit potential critiques and improvements to our system.
9. Conclusion
We presented BlendScape, a system that allows meeting participants to create blended meeting spaces that are tailored to their collaboration contexts and incorporate their own physical surroundings. Through implementing scenarios and conducting evaluations with 15 end-users, we demonstrated a rich set of distributed collaboration use cases that BlendScape can enable. Participants’ feedback was encouraging in that they could rapidly express their design intentions with BlendScape and envision using similar techniques to facilitate meetings in the future. Based on their feedback, we proposed improvements to BlendScape to mitigate challenges, e.g., incohesive or distracting scene elements. Future studies around BlendScape could explore how using BlendScape during live meetings impacts collaboration behavior and whether blended environments could mitigate challenges with traditional videoconferencing tools, e.g., fatigue and lack of co-presence.
Acknowledgements.
We would like to thank Amit Gulati and Henrik Turbell for their valuable feedback during the various stages of the project. We thank Sasa Junuzovic and Pat Sweeney for their valuable suggestions on working with NDI streams. Finally, we express our gratitude to study participants for their time.References
- (1)
- Agarwala et al. (2004) Aseem Agarwala, Mira Dontcheva, Maneesh Agrawala, Steven Drucker, Alex Colburn, Brian Curless, David Salesin, and Michael Cohen. 2004. Interactive digital photomontage. ACM Transactions on Graphics 23, 3 (Aug. 2004), 294–302. https://doi.org/10.1145/1015706.1015718
- Avidan and Shamir (2007) Shai Avidan and Ariel Shamir. 2007. Seam Carving for Content-Aware Image Resizing. In ACM SIGGRAPH 2007 Papers (SIGGRAPH ’07). Association for Computing Machinery, New York, NY, USA, 10–es. https://doi.org/10.1145/1275808.1276390
- Benabdallah et al. (2021) Gabrielle Benabdallah, Sam Bourgault, Nadya Peek, and Jennifer Jacobs. 2021. Remote Learners, Home Makers: How Digital Fabrication Was Taught Online During a Pandemic. In CHI ’21: CHI Conference on Human Factors in Computing Systems, Virtual Event / Yokohama, Japan, May 8-13, 2021, Yoshifumi Kitamura, Aaron Quigley, Katherine Isbister, Takeo Igarashi, Pernille Bjørn, and Steven Mark Drucker (Eds.). ACM, 350:1–350:14. https://doi.org/10.1145/3411764.3445450
- Brade et al. (2023) Stephen Brade, Bryan Wang, Mauricio Sousa, Sageev Oore, and Tovi Grossman. 2023. Promptify: Text-to-Image Generation through Interactive Prompt Exploration with Large Language Models. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, San Francisco , CA , USA. https://doi.org/10.48550/arXiv.2304.09337 arXiv:2304.09337 [cs]
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative Research in Psychology 3, 2 (2006), 77–101.
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models Are Few-Shot Learners. In Advances in Neural Information Processing Systems, Vol. 33. Curran Associates, Inc., 1877–1901.
- Buxton (2009) Bill Buxton. 2009. Mediaspace – Meaningspace – Meetingspace. Springer London, London, 217–231. https://doi.org/10.1007/978-1-84882-483-6_13
- Cai et al. (2021) Carrie J. Cai, Michelle Carney, Nida Zada, and Michael Terry. 2021. Breakdowns and Breakthroughs: Observing Musicians’ Responses to the COVID-19 Pandemic. In CHI ’21: CHI Conference on Human Factors in Computing Systems, Virtual Event / Yokohama, Japan, May 8-13, 2021, Yoshifumi Kitamura, Aaron Quigley, Katherine Isbister, Takeo Igarashi, Pernille Bjørn, and Steven Mark Drucker (Eds.). ACM, 571:1–571:13. https://doi.org/10.1145/3411764.3445192
- Chen et al. (2023) Xinyue Chen, Shuo Li, Shipeng Liu, Robin Fowler, and Xu Wang. 2023. MeetScript: Designing Transcript-based Interactions to Support Active Participation in Group Video Meetings. Proceedings of the ACM on Human-Computer Interaction abs/2309.12115 (2023). https://doi.org/10.48550/ARXIV.2309.12115 arXiv:2309.12115
- Choi and Diehm (2021) Jaz Hee-jeong Choi and Cade Diehm. 2021. Aesthetic flattening. Interactions 28, 4 (2021), 21–23. https://doi.org/10.1145/3468080
- Chung and Adar (2023a) John Joon Young Chung and Eytan Adar. 2023a. Artinter: AI-powered Boundary Objects for Commissioning Visual Arts. In Proceedings of the 2023 ACM Designing Interactive Systems Conference (DIS ’23). Association for Computing Machinery, New York, NY, USA, 1997–2018. https://doi.org/10.1145/3563657.3595961
- Chung and Adar (2023b) John Joon Young Chung and Eytan Adar. 2023b. PromptPaint: Steering Text-to-Image Generation Through Paint Medium-like Interactions. https://doi.org/10.1145/3586183.3606777 arXiv:2308.05184 [cs]
- Chung et al. (2022) John Joon Young Chung, Wooseok Kim, Kang Min Yoo, Hwaran Lee, Eytan Adar, and Minsuk Chang. 2022. TaleBrush: Sketching Stories with Generative Pretrained Language Models. In CHI Conference on Human Factors in Computing Systems. ACM, New Orleans LA USA, 1–19. https://doi.org/10.1145/3491102.3501819
- Coyne and Sproat (2001) Bob Coyne and Richard Sproat. 2001. WordsEye: An Automatic Text-to-Scene Conversion System. In Proceedings of the 28th Annual Conference on Computer Graphics and Interactive Techniques (SIGGRAPH ’01). Association for Computing Machinery, New York, NY, USA, 487–496. https://doi.org/10.1145/383259.383316
- Dang et al. (2023) Hai Dang, Frederik Brudy, George Fitzmaurice, and Fraser Anderson. 2023. WorldSmith: Iterative and Expressive Prompting for World Building with a Generative AI. arXiv:2308.13355 [cs]
- Dang et al. (2022) Hai Dang, Lukas Mecke, and Daniel Buschek. 2022. GANSlider: How Users Control Generative Models for Images Using Multiple Sliders with and without Feedforward Information. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3491102.3502141
- Dell et al. (2012) Nicola Dell, Vidya Vaidyanathan, Indrani Medhi, Edward Cutrell, and William Thies. 2012. ”Yours is better!”: participant response bias in HCI. In CHI Conference on Human Factors in Computing Systems, CHI ’12, Austin, TX, USA - May 05 - 10, 2012, Joseph A. Konstan, Ed H. Chi, and Kristina Höök (Eds.). ACM, 1321–1330. https://doi.org/10.1145/2207676.2208589
- El-Nouby et al. (2019) Alaaeldin El-Nouby, Shikhar Sharma, Hannes Schulz, R Devon Hjelm, Layla El Asri, Samira Ebrahimi Kahou, Yoshua Bengio, and Graham Taylor. 2019. Tell, Draw, and Repeat: Generating and Modifying Images Based on Continual Linguistic Instruction. In 2019 IEEE/CVF International Conference on Computer Vision (ICCV). IEEE, Seoul, Korea (South), 10303–10311. https://doi.org/10.1109/ICCV.2019.01040
- Fauville et al. (2021) G. Fauville, M. Luo, A.C.M. Queiroz, J.N. Bailenson, and J. Hancock. 2021. Zoom Exhaustion & Fatigue Scale. Computers in Human Behavior Reports 4 (2021), 100119. https://doi.org/10.1016/j.chbr.2021.100119
- Follmer et al. (2010) Sean Follmer, Hayes Raffle, Janet Go, Rafael Ballagas, and Hiroshi Ishii. 2010. Video Play: Playful Interactions in Video Conferencing for Long-Distance Families with Young Children. In Proceedings of the 9th International Conference on Interaction Design and Children. ACM, Barcelona Spain, 49–58. https://doi.org/10.1145/1810543.1810550
- Fuchsberger et al. (2021) Verena Fuchsberger, Janne Mascha Beuthel, Philippe Bentegeac, and Manfred Tscheligi. 2021. Grandparents and Grandchildren Meeting Online: The Role of Material Things in Remote Settings. In CHI ’21: CHI Conference on Human Factors in Computing Systems, Virtual Event / Yokohama, Japan, May 8-13, 2021, Yoshifumi Kitamura, Aaron Quigley, Katherine Isbister, Takeo Igarashi, Pernille Bjørn, and Steven Mark Drucker (Eds.). ACM, 478:1–478:14. https://doi.org/10.1145/3411764.3445191
- Gkioxari et al. (2020) Georgia Gkioxari, Jitendra Malik, and Justin Johnson. 2020. Mesh R-CNN. arXiv:1906.02739 [cs]
- Grønbæk et al. (2023a) Jens Emil Grønbæk, Wendy E Mackay, Marcel Borowski, Michel Beaudouin-Lafon, Eve Hoggan, and Clemens N Klokmose. 2023a. Mirrorverse: Live Tailoring of Video Conferencing Interfaces. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, San Francisco , CA , USA. https://doi.org/10.1145/3586183.3606767
- Grønbæk et al. (2021) Jens Emil Grønbæk, Banu Saatçi, Carla F. Griggio, and Clemens Nylandsted Klokmose. 2021. MirrorBlender: Supporting Hybrid Meetings with a Malleable Video-Conferencing System. In CHI ’21: CHI Conference on Human Factors in Computing Systems, Virtual Event / Yokohama, Japan, May 8-13, 2021, Yoshifumi Kitamura, Aaron Quigley, Katherine Isbister, Takeo Igarashi, Pernille Bjørn, and Steven Mark Drucker (Eds.). ACM, 451:1–451:13. https://doi.org/10.1145/3411764.3445698
- Grønbæk et al. (2023b) Jens Emil Sloth Grønbæk, Ken Pfeuffer, Eduardo Velloso, Morten Astrup, Melanie Isabel Sønderkær Pedersen, Martin Kjær, Germán Leiva, and Hans Gellersen. 2023b. Partially Blended Realities: Aligning Dissimilar Spaces for Distributed Mixed Reality Meetings. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI 2023, Hamburg, Germany, April 23-28, 2023, Albrecht Schmidt, Kaisa Väänänen, Tesh Goyal, Per Ola Kristensson, Anicia Peters, Stefanie Mueller, Julie R. Williamson, and Max L. Wilson (Eds.). ACM, 456:1–456:16. https://doi.org/10.1145/3544548.3581515
- Herskovitz et al. (2022) Jaylin Herskovitz, Yifei Cheng, Anhong Guo, Alanson P. Sample, and Michael Nebeling. 2022. XSpace: An Augmented Reality Toolkit for Enabling Spatially-Aware Distributed Collaboration. Proc. ACM Hum. Comput. Interact. 6, ISS (2022), 277–302. https://doi.org/10.1145/3567721
- Hu et al. (2023a) Erzhen Hu, Jens Emil Sloth Grønbæk, Austin Houck, and Seongkook Heo. 2023a. OpenMic: Utilizing Proxemic Metaphors for Conversational Floor Transitions in Multiparty Video Meetings. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI 2023, Hamburg, Germany, April 23-28, 2023, Albrecht Schmidt, Kaisa Väänänen, Tesh Goyal, Per Ola Kristensson, Anicia Peters, Stefanie Mueller, Julie R. Williamson, and Max L. Wilson (Eds.). ACM, 793:1–793:17. https://doi.org/10.1145/3544548.3581013
- Hu et al. (2023b) Erzhen Hu, Jens Emil Sloth Grønbæk, Wen Ying, Ruofei Du, and Seongkook Heo. 2023b. ThingShare: Ad-Hoc Digital Copies of Physical Objects for Sharing Things in Video Meetings. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI 2023, Hamburg, Germany, April 23-28, 2023, Albrecht Schmidt, Kaisa Väänänen, Tesh Goyal, Per Ola Kristensson, Anicia Peters, Stefanie Mueller, Julie R. Williamson, and Max L. Wilson (Eds.). ACM, 365:1–365:22. https://doi.org/10.1145/3544548.3581148
- Hunter et al. (2014) Seth E. Hunter, Pattie Maes, Anthony Tang, Kori M. Inkpen, and Susan M. Hessey. 2014. WaaZam!: supporting creative play at a distance in customized video environments. In CHI Conference on Human Factors in Computing Systems, CHI’14, Toronto, ON, Canada - April 26 - May 01, 2014, Matt Jones, Philippe A. Palanque, Albrecht Schmidt, and Tovi Grossman (Eds.). ACM, 1197–1206. https://doi.org/10.1145/2556288.2557382
- Hyrkas et al. (2023) Jeremy Hyrkas, Andrew D. Wilson, John Tang, Hannes Gamper, Hong Sodoma, Lev Tankelevitch, Kori Inkpen, Shreya Chappidi, and Brennan Jones. 2023. Spatialized Audio and Hybrid Video Conferencing: Where Should Voices be Positioned for People in the Room and Remote Headset Users?. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI 2023, Hamburg, Germany, April 23-28, 2023, Albrecht Schmidt, Kaisa Väänänen, Tesh Goyal, Per Ola Kristensson, Anicia Peters, Stefanie Mueller, Julie R. Williamson, and Max L. Wilson (Eds.). ACM, 794:1–794:14. https://doi.org/10.1145/3544548.3581085
- Jun and Nichol (2023) Heewoo Jun and Alex Nichol. 2023. Shap-E: Generating Conditional 3D Implicit Functions. https://doi.org/10.48550/arXiv.2305.02463 arXiv:2305.02463 [cs]
- Junuzovic et al. (2012) Sasa Junuzovic, Kori Inkpen, Tom Blank, and Anoop Gupta. 2012. IllumiShare: sharing any surface. In CHI Conference on Human Factors in Computing Systems, CHI ’12, Austin, TX, USA - May 05 - 10, 2012, Joseph A. Konstan, Ed H. Chi, and Kristina Höök (Eds.). ACM, 1919–1928. https://doi.org/10.1145/2207676.2208333
- Karnewar et al. (2023) Animesh Karnewar, Andrea Vedaldi, David Novotny, and Niloy J. Mitra. 2023. HOLODIFFUSION: Training a 3D Diffusion Model Using 2D Images. In 2023 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, Vancouver, BC, Canada, 18423–18433. https://doi.org/10.1109/CVPR52729.2023.01767
- Koh et al. (2022) Jing Yu Koh, Harsh Agrawal, Dhruv Batra, Richard Tucker, Austin Waters, Honglak Lee, Yinfei Yang, Jason Baldridge, and Peter Anderson. 2022. Simple and Effective Synthesis of Indoor 3D Scenes. https://doi.org/10.48550/arXiv.2204.02960 arXiv:2204.02960 [cs]
- Kumaravel et al. (2019) Balasaravanan Thoravi Kumaravel, Fraser Anderson, George W. Fitzmaurice, Bjoern Hartmann, and Tovi Grossman. 2019. Loki: Facilitating Remote Instruction of Physical Tasks Using Bi-Directional Mixed-Reality Telepresence. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology, UIST 2019, New Orleans, LA, USA, October 20-23, 2019, François Guimbretière, Michael S. Bernstein, and Katharina Reinecke (Eds.). ACM, 161–174. https://doi.org/10.1145/3332165.3347872
- Labrie et al. (2022) Audrey Labrie, Terrance Tin Hoi Mok, Anthony Tang, Michelle Lui, Lora Oehlberg, and Lev Poretski. 2022. Toward Video-Conferencing Tools for Hands-On Activities in Online Teaching. Proc. ACM Hum. Comput. Interact. 6, GROUP (2022), 10:1–10:22. https://doi.org/10.1145/3492829
- Ledo et al. (2018) David Ledo, Steven Houben, Jo Vermeulen, Nicolai Marquardt, Lora Oehlberg, and Saul Greenberg. 2018. Evaluation Strategies for HCI Toolkit Research. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems, CHI 2018, Montreal, QC, Canada, April 21-26, 2018. ACM, 36. https://doi.org/10.1145/3173574.3173610
- Lee et al. (2022) Minha Lee, Wonyoung Park, Sunok Lee, and Sangsu Lee. 2022. Distracting Moments in Videoconferencing: A Look Back at the Pandemic Period. In CHI ’22: CHI Conference on Human Factors in Computing Systems, New Orleans, LA, USA, 29 April 2022 - 5 May 2022, Simone D. J. Barbosa, Cliff Lampe, Caroline Appert, David A. Shamma, Steven Mark Drucker, Julie R. Williamson, and Koji Yatani (Eds.). ACM, 141:1–141:21. https://doi.org/10.1145/3491102.3517545
- Li et al. (2023) Yuheng Li, Haotian Liu, Qingyang Wu, Fangzhou Mu, Jianwei Yang, Jianfeng Gao, Chunyuan Li, and Yong Jae Lee. 2023. GLIGEN: Open-Set Grounded Text-to-Image Generation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 22511–22521.
- Liu et al. (2022) Vivian Liu, Han Qiao, and Lydia B. Chilton. 2022. Opal: Multimodal Image Generation for News Illustration. In The 35th Annual ACM Symposium on User Interface Software and Technology, UIST 2022, Bend, OR, USA, 29 October 2022 - 2 November 2022. ACM, 73:1–73:17. https://doi.org/10.1145/3526113.3545621
- Mirza and Osindero (2014) Mehdi Mirza and Simon Osindero. 2014. Conditional Generative Adversarial Nets. https://doi.org/10.48550/arXiv.1411.1784 arXiv:1411.1784 [cs, stat]
- Morikawa and Maesako (1998) Osamu Morikawa and Takanori Maesako. 1998. HyperMirror: Toward Pleasant-to-Use Video Mediated Communication System. In CSCW ’98, Proceedings of the ACM 1998 Conference on Computer Supported Cooperative Work, Seattle, WA, USA, November 14-18, 1998, Steven E. Poltrock and Jonathan Grudin (Eds.). ACM, 149–158. https://doi.org/10.1145/289444.289489
- Numan et al. (2023) Nels Numan, Daniele Giunchi, Benjamin Congdon, and Anthony Steed. 2023. Ubiq-Genie: Leveraging External Frameworks for Enhanced Social VR Experiences. In 2023 IEEE Conference on Virtual Reality and 3D User Interfaces Abstracts and Workshops (VRW). IEEE, Shanghai, China, 497–501. https://doi.org/10.1109/VRW58643.2023.00108
- O’Hara et al. (2011) Kenton O’Hara, Jesper Kjeldskov, and Jeni Paay. 2011. Blended interaction spaces for distributed team collaboration. ACM Trans. Comput. Hum. Interact. 18, 1 (2011), 3:1–3:28. https://doi.org/10.1145/1959022.1959025
- Olafenwa (2021) Ayoola Olafenwa. 2021. Simplifying Object Segmentation with PixelLib Library. (Jan. 2021).
- OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. https://doi.org/10.48550/arXiv.2303.08774 arXiv:2303.08774 [cs]
- Pejsa et al. (2016) Tomislav Pejsa, Julian Kantor, Hrvoje Benko, Eyal Ofek, and Andrew D. Wilson. 2016. Room2Room: Enabling Life-Size Telepresence in a Projected Augmented Reality Environment. In Proceedings of the 19th ACM Conference on Computer-Supported Cooperative Work & Social Computing, CSCW 2016, San Francisco, CA, USA, February 27 - March 2, 2016, Darren Gergle, Meredith Ringel Morris, Pernille Bjørn, and Joseph A. Konstan (Eds.). ACM, 1714–1723. https://doi.org/10.1145/2818048.2819965
- Radford et al. (2018) Alec Radford, Karthik Narasimhan, Tim Salimans, and Ilya Sutskever. 2018. Improving Language Understanding by Generative Pre-Training.
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Bjorn Ommer. 2022. High-Resolution Image Synthesis with Latent Diffusion Models. In 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, New Orleans, LA, USA, 10674–10685. https://doi.org/10.1109/CVPR52688.2022.01042
- Scupin (1997) Raymond Scupin. 1997. The KJ Method: A Technique for Analyzing Data Derived from Japanese Ethnology. Human Organization 56, 2 (1997), 233–237.
- Sharma et al. (2018) Shikhar Sharma, Dendi Suhubdy, Vincent Michalski, Samira Ebrahimi Kahou, and Yoshua Bengio. 2018. ChatPainter: Improving Text to Image Generation Using Dialogue. https://doi.org/10.48550/arXiv.1802.08216 arXiv:1802.08216 [cs]
- Sra et al. (2018) Misha Sra, Sergio Garrido-Jurado, and Pattie Maes. 2018. Oasis: Procedurally Generated Social Virtual Spaces from 3D Scanned Real Spaces. IEEE Transactions on Visualization and Computer Graphics 24, 12 (Dec. 2018), 3174–3187. https://doi.org/10.1109/TVCG.2017.2762691
- Tang et al. (2023) John Tang, Kori Inkpen, Sasa Junuzovic, Keri Mallari, Sean Rintel, Andrew Wilson, Shiraz Cupala, Tony Carbary, Abigail Sellen, and William Buxton. 2023. Perspectives: Creating Inclusive and Equitable Hybrid Meeting Experiences. Proceedings of the ACM on Human-Computer Interaction 7, CSCW2 (Oct. 2023).
- Tuddenham and Robinson (2009) Philip Tuddenham and Peter Robinson. 2009. Territorial coordination and workspace awareness in remote tabletop collaboration. In Proceedings of the 27th International Conference on Human Factors in Computing Systems, CHI 2009, Boston, MA, USA, April 4-9, 2009. ACM, 2139–2148. https://doi.org/10.1145/1518701.1519026
- Venolia et al. (2018) Gina Venolia, John C. Tang, Kori Inkpen, and Baris Unver. 2018. Wish you were here: being together through composite video and digital keepsakes. In Proceedings of the 20th International Conference on Human-Computer Interaction with Mobile Devices and Services, MobileHCI 2018, Barcelona, Spain, September 03-06, 2018, Lynne Baillie and Nuria Oliver (Eds.). ACM, 17:1–17:11. https://doi.org/10.1145/3229434.3229476
- Xia et al. (2018) Haijun Xia, Sebastian Herscher, Ken Perlin, and Daniel Wigdor. 2018. Spacetime: Enabling Fluid Individual and Collaborative Editing in Virtual Reality. In The 31st Annual ACM Symposium on User Interface Software and Technology, UIST 2018, Berlin, Germany, October 14-17, 2018. ACM, 853–866. https://doi.org/10.1145/3242587.3242597
- Yang et al. (2019) Jackie (Junrui) Yang, Christian Holz, Eyal Ofek, and Andrew D. Wilson. 2019. DreamWalker: Substituting Real-World Walking Experiences with a Virtual Reality. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19). Association for Computing Machinery, New York, NY, USA, 1093–1107. https://doi.org/10.1145/3332165.3347875
- Yu et al. (2019) Jiahui Yu, Zhe Lin, Jimei Yang, Xiaohui Shen, Xin Lu, and Thomas Huang. 2019. Free-Form Image Inpainting With Gated Convolution. In 2019 IEEE/CVF International Conference on Computer Vision (ICCV). IEEE, Seoul, Korea (South), 4470–4479. https://doi.org/10.1109/ICCV.2019.00457
- Yuan et al. (2021) Ye Yuan, Jan Cao, Ruotong Wang, and Svetlana Yarosh. 2021. Tabletop Games in the Age of Remote Collaboration: Design Opportunities for a Socially Connected Game Experience. In CHI ’21: CHI Conference on Human Factors in Computing Systems, Virtual Event / Yokohama, Japan, May 8-13, 2021, Yoshifumi Kitamura, Aaron Quigley, Katherine Isbister, Takeo Igarashi, Pernille Bjørn, and Steven Mark Drucker (Eds.). ACM, 436:1–436:14. https://doi.org/10.1145/3411764.3445512
- Zagermann et al. (2016) Johannes Zagermann, Ulrike Pfeil, Roman Rädle, Hans-Christian Jetter, Clemens Nylandsted Klokmose, and Harald Reiterer. 2016. When Tablets meet Tabletops: The Effect of Tabletop Size on Around-the-Table Collaboration with Personal Tablets. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems, San Jose, CA, USA, May 7-12, 2016. ACM, 5470–5481. https://doi.org/10.1145/2858036.2858224
- Zhang et al. (2023a) Lei Zhang, Ashutosh Agrawal, Steve Oney, and Anhong Guo. 2023a. VRGit: A Version Control System for Collaborative Content Creation in Virtual Reality. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, CHI 2023, Hamburg, Germany, April 23-28, 2023. ACM, 36:1–36:14. https://doi.org/10.1145/3544548.3581136
- Zhang et al. (2023b) Lvmin Zhang, Anyi Rao, and Maneesh Agrawala. 2023b. Adding Conditional Control to Text-to-Image Diffusion Models. https://doi.org/10.48550/arXiv.2302.05543 arXiv:2302.05543 [cs]
- Zhang et al. (2017) Richard Zhang, Jun-Yan Zhu, Phillip Isola, Xinyang Geng, Angela S. Lin, Tianhe Yu, and Alexei A. Efros. 2017. Real-Time User-Guided Image Colorization with Learned Deep Priors. ACM Transactions on Graphics 36, 4 (July 2017), 119:1–119:11. https://doi.org/10.1145/3072959.3073703
Appendix A Appendix
A.1. Additional Examples of Blended Environments
Overview of 12 examples of generated meeting environments, featuring prompts such as ”factory”, ”building treehouse”, ”birthday party” and ”wedding planning; traditional Chinese”.
A.2. GPT-aided Prompt Crafting for Image Generation
Crafting textual prompts that result in desired output with image generation models (e.g., Stable Diffusion) requires significant expertise, and trial and error (Brade et al., 2023). To ease the process of prompt creation for stable diffusion, BlendScape enhances users’ prompts by leveraging an LLM (in this case, GPT-3.5) to augment users’ prompts with additional keywords related to the Meeting Activity and Meeting Theme. The advantage of this approach is that it allows for the dynamic generation of keywords for image generation that suit the varying needs of users in the meeting.
For example, if the meeting activity is ”brainstorming”, and the theme of the meeting is ”mushroom forest”, then the system crafts a base prompt of Mushroom forest-themed environment for a brainstorming session. We then leverage single-shot prompting with GPT-3.5 to enhance the base prompt with contextually-relevant details:
System Prompt to GPT: Your task is to help the user create a Stable Diffusion prompt to generate an environment design. The user will specify an activity to occur in the environment and/or a theme for the space. You will provide a list of 4-5 types of objects to put in the environment and 4-5 distinct characteristics that describe the environment. The characteristics must be detailed and designed to generate visually appealing and cohesive results. Here is an example for a brainstorming activity:
Objects: "whiteboards, plants, chairs, small tables",
Environment Characteristics: "bright, open space, natural light, refreshing atmosphere, varied textures"
Subsequently, depending on the base prompt derived from the user’s input prompt, BlendScape crafts an input to GPT-3.5:
Input Prompt (constructed by BlendScape) to GPT-3.5: Provide a list of 4-5 types of objects to put in this environment and 4-5 characteristics that describe this environment: Mushroom forest-themed environment for a brainstorming session. Return the output as comma-separated strings in JSON format: {Objects: string, Environment Characteristics: string}
The results of such a prompting strategy to GPT results in the following output.
Objects: "Mushrooms, Trees, Fairies, Moss-covered rocks, Lanterns",
Environment Characteristics": "Enchanting, Magical, Misty, Whimsical, Serene"
Additionally, BlendScape adds a fixed set of terms - “highly detailed, intricate, sharp focus, smooth” that we found to improve the results. In this case, the final prompt provided to Stable Diffusion is:
Mushroom forest-themed environment for a brainstorming session; Giant mushrooms, Fairy houses, Moss-covered rocks, Glowing mushrooms, Enchanted flowers; Enchanting, Magical, Misty, Whimsical, Serene; highly detailed, intricate, sharp focus, smooth
Fig. 11 from our Storytelling with Family scenario shows the result of an image generation with such a prompt.
A.3. User Study Scenarios
Overview of the ”Vacation Planning Scenario”. Two friends use BlendScape to get into the spirit of planning a trip to Paris. In Task 1 of our user study, we asked participants to compare three pre-canned environments that preserve their video backgrounds to different extents. This figure shows the three generated environments, with the first environment preserving the video backgrounds the most, and the third environment preserving the video backgrounds the least.
Overview of the ”Game Stream” Scenario: A gamer incorporates his Minecraft stream into a blended environment to make his viewers feel more connected to the gameplay. Participants compared four versions of spaces generated via image-to-image techniques, using two different priors and varying levels of background preservation. This figure shows the four generated environments, of which two preserve the video backgrounds and two do not. Each environment has a unique placement of users and the Minecraft stream.